The salt copper(II) sulfate is soluble in water. When 2.27 g of CuSO4 is dissolved in 107.00 g of water, the temperature of the solution increases from 25.00 to 27.42 °C. Based on this observation, calculate the enthalpy of dissolution of CuSO4 (in kJ/mol).

Assume that the specific heat of the solution is 4.184 J/g °C and that the heat absorbed by the calorimeter is negligible.

ΔHdissolution = ______ kJ/mol
1
Expert's answer
2019-04-08T07:56:28-0400

Find the amount of heat released after dissolving:

Q = cm(T2 -- T1), where c -- specific heat of the solution; m -- mass of the solution;


"Q = 4.184 * (107+2.27) * (27.42-25) = 1106.4 (J)"

We've found the amount of heat released after dissolving of 2.27 g of CuSO4, find the heat released after dissolving of 1 mole (159.6 g):


"H(dis) = 1106.4*159.6\/2.27 = 77789 (J\/mol) = 77.79 (kJ\/mol)"

Answer

77.79 kJ/mol is the enthalpy of dissolution of CuSO4.